### v1

@@ 2.3 Technical lemmas
In this part, we provide three technical lemmas for our discussions. The first one is about the polar cones.
Lemma 2.2 ([40, Corollary 11.25(d)]).
Let ${\mathcal{H}}:{\mathcal{E}}\to{\mathcal{F}}$ be a linear operator, and $K\subset{\mathcal{F}}$ be a nonempty closed convex cone. Then the set $C:=\{z\in{\mathcal{E}}\mid{\mathcal{A}}z\in K\}$ is also a closed convex cone and one has
$C^{\circ}={\rm cl}\,\{{\mathcal{A}}^{*}v\mid v\in K^{\circ}\},$
@@
The following lemma gives a variational characterization of self-adjoint positive definite operators. One can also see [18, Proposition 3.1] for a more straightforward proof based on the Moreau decomposition.
Lemma 2.3 ([17, Theorem 3.6]).
@@
$\langle z,{\mathcal{H}}z\rangle>0\quad\forall z\in{\mathcal{E}}\quad\Leftrightarrow\quad\left\{\begin{array}[]{ll}\langle z,{\mathcal{H}}z\rangle>0\ \forall z\in C\backslash\{0\}\ \mbox{and}\\[2.84526pt] \langle z,{\mathcal{H}}^{-1}z\rangle>0\ \forall z\in C^{\circ}\backslash\{0\}.\end{array}\right.$
The following lemma discusses Robinson’s constraint qualification [33], via its equivalent reformulation [6, Proposition 2.97], of a constraint system.
@@
Let ${\mathcal{H}}:{\mathcal{E}}\to{\mathcal{F}}$ be a linear operator, and $K\subseteq{\mathcal{F}}$ be a nonempty closed convex cone. Then, Robinson’s constraint qualification holds at any feasible point
$(\bar{z};\bar{v})\in\Big{\{}(z;v)\in{\mathcal{E}}\times{\mathcal{F}}\,\big{|}\,z-{\mathcal{H}}^{*}v=0,\,\frac{1}{2}(\|v\|^{2}-1)=0,\,v\in K\Big{\}},$
in the sense that for any given $(\Delta_{z};\delta;\Delta_{v})\in{\mathcal{E}}\times{\mathds{R}}\times{\mathcal{F}}$, one can find $z\in{\mathcal{E}}$, $v\in{\mathcal{F}}$ and $u\in{\mathcal{T}}_{K}(\bar{v})$ such that
@@
Proof.
Note that ${\mathcal{T}}_{K}(\bar{v})$ is a closed convex cone. From [40, Example 6.24] we know that ${\mathcal{N}}_{K}(\bar{v})=({\mathcal{T}}_{K}(\bar{v}))^{\circ}$. Then by Moreau’s decomposition theorem [37, Theorem 31.5] we know that one can uniquely decompose $\Delta_{v}$ by
$\Delta_{v}=\Delta_{v}^{\prime}+\Delta_{v}^{\prime\prime},\quad\Delta_{v}^{\prime}\in{\mathcal{T}}_{K}(\bar{v}),\quad\Delta_{v}^{\prime\prime}\in{\mathcal{N}}_{K}(\bar{v}),\quad\langle\Delta_{v}^{\prime},\Delta_{v}^{\prime\prime}\rangle=0.$
One can take $v_{0}:=\Delta_{v}^{\prime\prime}\in{\mathcal{N}}_{K}(\bar{v})$ and $u_{0}=\Delta_{v}^{\prime}\in{\mathcal{T}}_{K}(\bar{v})$, so that $v_{0}+u_{0}=\Delta_{v}$. Moreover, it is easy to see that
@@
Then, by letting $v:=v_{0}-\lambda\bar{v}$, one has $u:=u_{0}+\lambda\bar{v}\in{\mathcal{T}}_{K}(\bar{z})$ and $z:=\Delta_{z}+{\mathcal{H}}^{*}v$. Consequently, one has (2.19a) and (2.19c) hold. Moreover, since $\|\bar{v}\|=1$, it holds that $\langle\bar{v},v\rangle=\langle\bar{v},v_{0}-\lambda\bar{v}\rangle=\langle\bar{v},v_{0}\rangle-\lambda=\delta$, so that (2.19b) holds. This completes the proof. ∎
@@ 3 Implications of the Aubin property for NLSDP
In this section, we study the implications of the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) for the NLSDP problem (1.8). Throughout this section, we set $G(x):=(h(x);g(x))=(h(x);g_{1}(x);\ldots;g_{J}(x))$, ${\mathcal{Y}}:={\mathds{R}}^{m}\times\prod_{j=1}^{J}{\mathcal{S}}^{p_{j}}$ and ${\mathcal{K}}:=\{0\in{\mathds{R}}^{m}\}\times\prod_{j=1}^{J}{\mathcal{S}}_{+}^{p_{j}}$ in (1.1), and make no distinction between (1.1) and (1.8).
@@ 3.1 A reduction method for NLSDP
This part exploits the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ in (1.6). By using a reduction method, we can get more accessible results from the original formula of the Mordukhovich criterion.
Note that if ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) has the Aubin property at $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})$, the mapping ${\mathbb{S}}_{\rm GE}$, defined by
@@
also has the Aubin property at $0\in{\mathcal{X}}$ for $\bar{x}$. Moreover, according to the Mordukhovich criterion [27], the latter Aubin property holds if and only if ${\mathcal{D}}^{*}{{\mathbb{S}}}_{\rm GE}(0,\bar{x})(0)=\{0\}$. Since ${\mathcal{K}}$ is $C^{2}$-cone reducible at every $y\in{\mathcal{K}}$, one can utilize the second-order chain rule developed in [31, Theorem 7], as a generalization of [29, Theorem 3.4], to get the following result.
@@
Let $\bar{x}$ be a locally optimal solution of problem (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$. Suppose that $\bar{x}$ is nondegenerate, i.e., (2.14) holds. Then one has
@@
The proof follows from [31, Theorem 20], in which ${\mathcal{K}}$ was assumed to be a second-order cone, but it still holds when ${\mathcal{K}}$ is a closed convex set. ∎
@@
Let $\bar{x}$ be a nondegenerate locally optimal solution of problem (1.8) such that $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\ldots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$. For $j\in\{1,\ldots,J\}$, define $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and write its eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ as in (2.4) with $P_{j}=\big{(}(P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}}\big{)}$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Then, the mapping ${\mathbb{S}}_{\rm GE}$ defined by (3.1) has the Aubin property at 00 for $\bar{x}$ if and only if
$\begin{array}[]{ll}\displaystyle{\mathcal{Q}}d\notin{\sf rge}(\nabla h(\bar{x}))-\Bigg{\{}\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}\cup\gamma_{j}}\begin{pmatrix}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}&\widetilde{U}^{j}_{\beta_{j}\gamma_{j}}\\[5.69054pt] ({\widetilde{U}}^{j}_{\beta_{j}\gamma_{j}})^{\top}&\widetilde{U}^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}(P_{j})_{\beta_{j}\cup\gamma_{j}}^{\top}\\[17.07164pt] \quad\bigg{|}\ \begin{array}[]{ll}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)\big{(}(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\beta_{j}}\big{)}\\[8.53581pt] \forall{\widetilde{U}^{j}}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|},{\widetilde{U}^{j}}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|},\ j=1,\ldots,J\end{array}\Bigg{\}}\quad\forall d\in{\sf aff}({\mathcal{C}}(\bar{x}))\backslash\{0\},\end{array}$ (3.2)
where ${\mathcal{Q}}$ is the linear operator defined in (2.16).
@@
On the one hand, since $h(\bar{x})=0$, it is easy to see from the definition of the coderivative in (2.3) that
@@
On the other hand, for all $j=1,\ldots,J$, by applying Lemma 2.1 to $A_{j}=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and $V^{j}:=g^{\prime}_{j}(\bar{x})d$ we know that $U^{j}\in D^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|p_{j}|}}(g_{j}(\bar{x}),\bar{\Gamma}_{j})(g^{\prime}_{j}(\bar{x})d)$ holds if and only if $U:=U^{j}$ and $V:=g^{\prime}_{j}(\bar{x})d$ can be expressed as in (2.8) with $\alpha:=\alpha_{j}$, $\beta:=\beta_{j}$ and $\gamma:=\gamma_{j}$ such that (2.9) holds. Since $\bar{x}$ is nondegenerate, from the formulation of ${\sf aff}({\mathcal{C}}(\bar{x}))$ given in (2.15) one can see that (3.3) holds if and only if
@@
where each $\lambda^{j}_{i}$ is the $i$-th diagonal element of $\Lambda^{j}$. Note that for any $d\in{\mathcal{X}}$ one has $g^{\prime}_{j}(\bar{x})d\in{\mathcal{S}}^{p_{j}}$ and $(g_{j}(\bar{x}))^{\dagger}\in{\mathcal{S}}^{p_{j}}$. Therefore, it holds that
$\begin{array}[]{l}2\big{\langle}\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\big{\rangle}\\[5.69054pt] =\big{\langle}\bar{\Gamma}_{j}(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger},g^{\prime}_{j}(\bar{x})d\big{\rangle}+\big{\langle}(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\bar{\Gamma}_{j},g^{\prime}_{j}(\bar{x})d\big{\rangle},\\[5.69054pt] =\big{\langle}\nabla g_{j}(\bar{x})(\bar{\Gamma}_{j}(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}+(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\bar{\Gamma}_{j}),d\big{\rangle}\quad\forall j=1,\dots,J.\end{array}$ (3.6)
Moreover, for any $d\in{\mathcal{X}}$ and $j\in\{1,\ldots,J\}$, one can get from the eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ that
$\begin{array}[]{ll}(g_{j}(\bar{x}))^{\dagger}(g_{j}^{\prime}(\bar{x})d)\bar{\Gamma}_{j}&=P_{j}\begin{pmatrix}(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\\ &0\\ &&0\end{pmatrix}P_{j}^{\top}(g_{j}^{\prime}(\bar{x})d)P_{j}\begin{pmatrix}0\\ &0\\ &&\Lambda^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}P_{j}^{\top}\\[19.91692pt] &=P_{j}\begin{pmatrix}0&0&(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}\Lambda^{j}_{\gamma_{j}\gamma_{j}}\\ 0&0&0\\ 0&0&0\end{pmatrix}P_{j}^{\top}.\end{array}$ (3.7)
Similarly, one also has that
$\bar{\Gamma}_{j}(g_{j}^{\prime}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}=P_{j}\begin{pmatrix}0&0&0\\ 0&0&0\\ \Lambda^{j}_{\gamma_{j}\gamma_{j}}(\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}})^{\top}(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}&0&0\end{pmatrix}P_{j}^{\top}.$ (3.8)
Then, by putting (3.6), (3.7) and (3.8) together, one can see that for any $d\in{\mathcal{X}}$ and any $j=1,\ldots,J$, it holds that
$\begin{array}[]{l}2\langle\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\rangle\\[5.69054pt] =\left\langle\nabla g_{j}(\bar{x})P_{j}\begin{pmatrix}0&0&(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}\Lambda^{j}_{\gamma_{j}\gamma_{j}}\\ 0&0&0\\ \Lambda^{j}_{\gamma_{j}\gamma_{j}}(\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}})^{\top}(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}&0&0\end{pmatrix}P_{j}^{\top},d\right\rangle.\end{array}$ (3.9)
Note that for any $U^{j}\in\mathds{U}^{j}_{d}$, $j=1,\ldots,J$, it holds that
@@
and $\Sigma^{j}_{\alpha_{j}\gamma_{j}}\circ\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}=(E_{\alpha_{j}\gamma_{j}}-\Sigma_{\alpha_{j}\gamma_{j}})\circ\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}$. Then, from the definition of $\Sigma_{\alpha_{j}\gamma_{j}}$ in (3.5) one can see that for any $i\in\alpha_{j}$ and $k\in\gamma_{j}$ it holds that
@@
Consequently, one has that ${\lambda^{j}_{i}}\widetilde{U}^{j}_{ik}+{\lambda^{j}_{k}}\widetilde{V}^{j}_{ik}=0$. Therefore, it holds that
@@
$(\widetilde{U}^{j})_{\alpha_{j}\gamma_{j}}+(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}\Lambda^{j}_{\gamma_{j}\gamma_{j}}=0,$
@@
which, together with (3.9) and (2.16), implies that (3.2) holds if and only if (3.4) holds. This completes the proof. ∎
@@
When the assumptions in Proposition 3.1 hold, one can define the linear operator ${\mathcal{B}}:{\mathcal{X}}\to{\mathds{R}}^{m}\times\prod_{j=1}^{J}({\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|}\times{\mathcal{S}}^{|\gamma_{j}|})$ by

### v2

@@ 2.3 Technical lemmas
In this part, we provide three technical lemmas for our discussions. These results are not specialized only for the problems considered here. The first lemma is about the polar cone.
Lemma 2.3 ([43, Corollary 11.25(d)]).
Let ${\mathcal{H}}:{\mathcal{E}}\to{\mathcal{F}}$ be a linear operator and $K\subseteq{\mathcal{F}}$ be a closed nonempty convex cone. Then the set $C:=\{z\in{\mathcal{E}}\mid{\mathcal{H}}z\in K\}$ is also a closed convex cone and one has
$C^{\circ}={\rm cl}\,\{{{\mathcal{H}}}^{*}v\mid v\in K^{\circ}\},$
@@
The following lemma gives a variational characterization of self-adjoint positive definite operators. One can also see [20, Proposition 3.1] for a more straightforward proof based on the Moreau decomposition.
Lemma 2.4 ([18, Theorem 3.6]).
@@
$\langle z,{\mathcal{H}}z\rangle>0\quad\forall z\in{\mathcal{E}}\setminus\{0\}\quad\Leftrightarrow\quad\left\{\begin{array}[]{ll}\langle z,{\mathcal{H}}z\rangle>0\ \forall z\in C\backslash\{0\}\ \mbox{and}\\[2.84526pt] \langle z,{\mathcal{H}}^{-1}z\rangle>0\ \forall z\in C^{\circ}\backslash\{0\}.\end{array}\right.$
The following lemma discusses Robinson’s constraint qualification [36] of a special constraint system. Recall that for the general constraint system $G(x)\in{\mathcal{K}}$ in (1.1), we say that Robinson’s constraint qualification holds at a point $x\in{\mathcal{X}}$ such that $G(x)\in{\mathcal{K}}$ if $0\in{{\sf int}\,}\{G(x)+G^{\prime}(x){\mathcal{X}}-{\mathcal{K}}\}$. According to [7, Proposition 2.97 & Corollary 2.98], this constraint qualification is equivalent to
$G^{\prime}(x){\mathcal{X}}+{\mathcal{T}}_{{\mathcal{K}}}(G(x))={\mathcal{Y}}.$
Moreover, if Robinson’s constraint qualification holds at a locally optimal solution $\bar{x}$ of (1.1), then the set of multipliers ${\mathcal{M}}(\bar{x})$ is non-empty, convex and compact [7, Theorem 3.9].
@@
Let ${\mathcal{H}}:{\mathcal{E}}\to{\mathcal{F}}$ be a linear operator and $K\subseteq{\mathcal{F}}$ be a nonempty closed convex cone. Then, Robinson’s constraint qualification holds at any feasible point to the constraint system
$z-{\mathcal{H}}^{*}v=0,\quad\frac{1}{2}(\|v\|^{2}-1)=0,\quad\mbox{and}\quad v\in K.$ (2.19)
Proof.
Suppose that $(\bar{z};\bar{v})\in{\mathcal{E}}\times{\mathcal{F}}$ is a given feasible point to the constraint system (2.19). Let $(\Delta_{z};\delta;\Delta_{v})\in{\mathcal{E}}\times{\mathds{R}}\times{\mathcal{F}}$ be arbitrarily chosen. To show that Robinson’s constraint qualification holds at $(\bar{z};\bar{v})$ we need to find $z\in{\mathcal{E}}$, $v\in{\mathcal{F}}$ and $u\in{\mathcal{T}}_{K}(\bar{v})$ such that
@@
Since $K$ is a closed convex cone, from [43, Example 6.24] we know that ${\mathcal{N}}_{K}(\bar{v})=({\mathcal{T}}_{K}(\bar{v}))^{\circ}$. Then by Moreau’s decomposition theorem [40, Theorem 31.5] we know that one can uniquely decompose $\Delta_{v}$ by $\Delta_{v}=\Delta_{v}^{\prime}+\Delta_{v}^{\prime\prime}$ such that
$\Delta_{v}^{\prime}\in{\mathcal{T}}_{K}(\bar{v}),\quad\Delta_{v}^{\prime\prime}\in{\mathcal{N}}_{K}(\bar{v}),\quad\mbox{and}\quad\langle\Delta_{v}^{\prime},\Delta_{v}^{\prime\prime}\rangle=0.$
One can take $v_{0}:=\Delta_{v}^{\prime\prime}\in{\mathcal{N}}_{K}(\bar{v})$ and $u_{0}:=\Delta_{v}^{\prime}\in{\mathcal{T}}_{K}(\bar{v})$, so that $v_{0}+u_{0}=\Delta_{v}$. Moreover, it is easy to see from $\bar{v}\in K$ that
@@
Then, by letting $v:=v_{0}-\lambda\bar{v}$ and $u:=u_{0}+\lambda\bar{v}$ one has $u\in{\mathcal{T}}_{K}(\bar{v})$ and that (2.20c) holds. Moreover, since $\|\bar{v}\|=1$, we have $\langle\bar{v},v\rangle=\langle\bar{v},v_{0}-\lambda\bar{v}\rangle=\langle\bar{v},v_{0}\rangle-\lambda=\delta$, so that (2.20b) is valid. Then, letting $z:=\Delta_{z}+{\mathcal{H}}^{*}v$, we have that (2.20a) holds. This completes the proof. ∎
@@ 3 Implications of the Aubin property for NLSDP
In this section, we study the implications of the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) for the NLSDP problem (1.8). Throughout this section, we set $G(x):=(h(x);g(x))=(h(x);g_{1}(x);\ldots;g_{J}(x))$, ${\mathcal{Y}}:={\mathds{R}}^{m}\times\prod_{j=1}^{J}{\mathcal{S}}^{p_{j}}$ and ${\mathcal{K}}:=\{0\in{\mathds{R}}^{m}\}\times\prod_{j=1}^{J}{\mathcal{S}}_{+}^{p_{j}}$ in (1.1). Moreover, we make no distinction between the general constraint optimization problem (1.1) and the NLSDP problem (1.8) for convenience.
@@ 3.1 A reduction method for NLSDP
This part exploits the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ in (1.6). Note that if ${\mathbb{S}}_{{\rm KKT}}$ in (1.6) has the Aubin property at $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})$, the tilt perturbation solution mapping ${\mathbb{S}}_{\rm GE}$, defined by
@@
also has the Aubin property at $0\in{\mathcal{X}}$ for $\bar{x}$. Moreover, according to the Mordukhovich criterion [29], the latter Aubin property holds if and only if
${\mathcal{D}}^{*}{{\mathbb{S}}}_{\rm GE}(0,\bar{x})(0)=\{0\}.$
Next, we use a reduction method to get more accessible results from the original formula of the Mordukhovich criterion. Since ${\mathcal{K}}$ is $C^{2}$-cone reducible at every $y\in{\mathcal{K}}$, one can utilize the second-order chain rule developed in [33, Theorem 7], as a generalization of [31, Theorem 3.4], to get the following result.
@@
Let $\bar{x}$ be a stationary point of (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$. Suppose that $\bar{x}$ is nondegenerate, i.e., (2.14) holds. Then one has
@@
The proof follows directly from [33, Theorem 20], in which ${\mathcal{K}}$ was assumed to be a second-order cone, but it still holds when ${\mathcal{K}}$ is a general closed convex set. ∎
@@
Let $\bar{x}$ be a nondegenerate stationary point of (1.8) such that $\bar{y}=(\bar{\zeta};\bar{\Gamma}_{1};\ldots;\bar{\Gamma}_{J})\in{\mathcal{M}}(\bar{x})$. For $j\in\{1,\ldots,J\}$, define $A_{j}:=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and write its eigenvalue decomposition $A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ as in (2.4) with $P_{j}=\big{(}(P_{j})_{\alpha_{j}},(P_{j})_{\beta_{j}},(P_{j})_{\gamma_{j}}\big{)}$ being the corresponding orthogonal matrix and $\Lambda^{j}$ being the corresponding diagonal matrix of eigenvalues. Then, the mapping ${\mathbb{S}}_{\rm GE}$ defined by (3.1) has the Aubin property at 00 for $\bar{x}$ if and only if
$\begin{array}[]{ll}\displaystyle{\mathcal{Q}}d\notin{\sf rge}(\nabla h(\bar{x}))-\Bigg{\{}\sum_{j=1}^{J}\nabla g_{j}(\bar{x})(P_{j})_{\beta_{j}\cup\gamma_{j}}\begin{pmatrix}\widetilde{U}_{\beta_{j}\beta_{j}}^{j}&\widetilde{U}^{j}_{\beta_{j}\gamma_{j}}\\[5.69054pt] ({\widetilde{U}}^{j}_{\beta_{j}\gamma_{j}})^{\top}&\widetilde{U}^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}(P_{j})_{\beta_{j}\cup\gamma_{j}}^{\top}\\[17.07164pt] \qquad\bigg{|}\ \begin{array}[]{ll}\widetilde{U}^{j}_{\beta_{j}\beta_{j}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{|\beta_{j}|}}(0,0)\big{(}(P_{j})_{\beta_{j}}^{\top}(g^{\prime}_{j}(\bar{x})d)(P_{j})_{\beta_{j}}\big{)}\\[8.53581pt] \forall{\widetilde{U}^{j}}_{\beta_{j}\gamma_{j}}\in{\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|},{\widetilde{U}^{j}}_{\gamma_{j}\gamma_{j}}\in{\mathcal{S}}^{|\gamma_{j}|},\ j=1,\ldots,J\end{array}\Bigg{\}}\quad\forall d\in{\sf aff}({\mathcal{C}}(\bar{x}))\backslash\{0\},\end{array}$ (3.2)
where ${\mathcal{Q}}$ is the linear operator defined by (2.16).
@@
In the following, we will reformulate the right-hand side of (3.3) into an explicit form by using specific formulas of coderivatives for normal cones. On the one hand, since $h(\bar{x})=0$, it is easy to see from the definition of the coderivative in (2.3) that
@@
On the other hand, for all $j=1,\ldots,J$, by applying Lemma 2.1 to $A:=A_{j}=g_{j}(\bar{x})+\bar{\Gamma}_{j}$ and $V:=V^{j}:=g^{\prime}_{j}(\bar{x})d$ we know that $U^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{S}}_{+}^{p_{j}}}(g_{j}(\bar{x}),\bar{\Gamma}_{j})(g^{\prime}_{j}(\bar{x})d)$ holds if and only if $U:=U^{j}$ and $V$ can be expressed as in (2.8) with $\alpha:=\alpha_{j}$, $\beta:=\beta_{j}$ and $\gamma:=\gamma_{j}$ such that (2.9) holds. Since $\bar{x}$ is nondegenerate, from the formula of ${\sf aff}({\mathcal{C}}(\bar{x}))$ given in (2.15) one can see that (3.3) holds if and only if
@@
where $/0$ 0 is defined to be $1$ and each $\lambda^{j}_{i}$ is the $i$-th diagonal element of $\Lambda^{j}$. Note that for any $U^{j}\in\mathds{U}^{j}_{d}$, $j=1,\ldots,J$, it holds that
@@
and $\Sigma^{j}_{\alpha_{j}\gamma_{j}}\circ\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}=(E_{\alpha_{j}\gamma_{j}}-\Sigma_{\alpha_{j}\gamma_{j}})\circ\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}$. Then, from the definition of $\Sigma^{j}_{\alpha_{j}\gamma_{j}}$ in (3.5) one can see that for any $i\in\alpha_{j}$ and $k\in\gamma_{j}$ it holds that
@@
Consequently, one has ${\lambda^{j}_{i}}\widetilde{U}^{j}_{ik}+{\lambda^{j}_{k}}\widetilde{V}^{j}_{ik}=0$. Therefore, it holds that
$\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}+(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}\Lambda^{j}_{\gamma_{j}\gamma_{j}}=0,\quad j=1,\ldots,J.$ (3.6)
Next, we provide an equivalent reformulation of the last term [^1] This term was known as the “sigma term” (cf. [5, p. 177]). in (2.16) for defining the linear operator ${\mathcal{Q}}$, for the purpose of studying the left-hand side of (3.2). Note that for any $d\in{\mathcal{X}}$ one has $g^{\prime}_{j}(\bar{x})d\in{\mathcal{S}}^{p_{j}}$ and $(g_{j}(\bar{x}))^{\dagger}\in{\mathcal{S}}^{p_{j}}$. Therefore, it holds that
$\begin{array}[]{l}2\big{\langle}\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\big{\rangle}\\[5.69054pt] =\big{\langle}\bar{\Gamma}_{j}(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger},g^{\prime}_{j}(\bar{x})d\big{\rangle}+\big{\langle}(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\bar{\Gamma}_{j},g^{\prime}_{j}(\bar{x})d\big{\rangle},\\[5.69054pt] =\big{\langle}\nabla g_{j}(\bar{x})\big{(}\bar{\Gamma}_{j}(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}+(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\bar{\Gamma}_{j}\big{)},d\big{\rangle}\quad\forall j=1,\dots,J.\end{array}$ (3.7)
@@
One can see from the eigenvalue decomposition of $g_{j}(\bar{x})+\bar{\Gamma}_{j}=A_{j}=P_{j}\Lambda^{j}P_{j}^{\top}$ and $\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}=(P_{j})_{\alpha_{j}}^{\top}(g_{j}^{\prime}(\bar{x})d)(P_{j})_{\gamma_{j}}$ that
$\begin{array}[]{ll}(g_{j}(\bar{x}))^{\dagger}(g_{j}^{\prime}(\bar{x})d)\bar{\Gamma}_{j}=P_{j}\begin{pmatrix}(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\\ &0\\ &&0\end{pmatrix}P_{j}^{\top}(g_{j}^{\prime}(\bar{x})d)P_{j}\begin{pmatrix}0\\ &0\\ &&\Lambda^{j}_{\gamma_{j}\gamma_{j}}\end{pmatrix}P_{j}^{\top}\\[19.91692pt] \qquad=P_{j}\begin{pmatrix}0&0&(\Lambda^{j}_{\alpha_{j}\alpha_{j}})^{-1}\widetilde{V}^{j}_{\alpha_{j}\gamma_{j}}\Lambda^{j}_{\gamma_{j}\gamma_{j}}\\ 0&0&0\\ 0&0&0\end{pmatrix}P_{j}^{\top}=P_{j}\begin{pmatrix}0&0&-\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}\\ 0&0&0\\ 0&0&0\end{pmatrix}P_{j}^{\top},\end{array}$ (3.8)
where the last equality comes from (3.6). Similarly, one also has that
$\bar{\Gamma}_{j}(g_{j}^{\prime}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}=P_{j}\begin{pmatrix}0&0&0\\ 0&0&0\\ -\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}&0&0\end{pmatrix}P_{j}^{\top}.$ (3.9)
Then, by putting (3.7), (3.8) and (3.9) together, one can see that
$\begin{array}[]{l}2\langle\bar{\Gamma}_{j},(g^{\prime}_{j}(\bar{x})d)(g_{j}(\bar{x}))^{\dagger}(g^{\prime}_{j}(\bar{x})d)\rangle=\left\langle\nabla g_{j}(\bar{x})P_{j}\begin{pmatrix}0&0&-\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}\\ 0&0&0\\ -\widetilde{U}^{j}_{\alpha_{j}\gamma_{j}}&0&0\end{pmatrix}P_{j}^{\top},d\right\rangle.\end{array}$ (3.10)
@@
Now, one can see from the definition of ${\mathcal{Q}}$ in (2.16) and (3.10) that (3.2) holds if and only if (3.4) holds. This completes the proof. ∎
@@
Proposition 3.1 may seem complicated and confusing, but in fact, it represents the first step of our reduction method. Taking a look at (3.2), only the $\beta_{j}$ and $\gamma_{j}$ parts remain, and we have reduced the $\alpha_{j}$ part in $g_{j}(\bar{x})+\bar{\Gamma}_{j}$. When the assumptions in Proposition 3.1 hold, one can further define the linear operator ${\mathcal{B}}:{\mathcal{X}}\to{\mathds{R}}^{m}\times\prod_{j=1}^{J}({\mathds{R}}^{|\beta_{j}|\times|\gamma_{j}|}\times{\mathcal{S}}^{|\gamma_{j}|})$ by